DIMACS

Center for Discrete Mathematics & Theoretical Computer Science

DIMACS EDUCATIONAL MODULES SERIES

MODULE 03-2 Facility Location Problems

Date prepared: March 27,2003

R. S. (Chuck) Tiberio ^{1} Wellesley High School Wellesley, MA 02481 email: tiberio@tiac.net

DIMACS Center, CoRE Bldg., Rutgers University, 96 Frelinghuysen Road, Piscataway, NJ 08854-8018

TEL: 732-445-5928 • FAX: 732-445-5932 • EMAIL: center@dimacs.rutgers.edu

                       Web: http://dimacs.rutgers.edu/

Founded as a National Science Foundation Science and Technology Center and a Joint Project of Rutgers University,

Princeton University, AT&T Labs - Research, Bell Labs, NEC Laboratories America and Telcordia Technologies with affiliated members Avaya Labs, HP Labs, IBM Research, Microsoft Research.

1Written under the auspices of DREI98, supported by NSF Grant DMS 9412914

Module Description Information

• Title:

Facility Location Problems

• Author(s):

R. S. (Chuck) Tiberio

• Abstract:

This module investigates the placement of emergency and service facilities in a network of towns connected by roads. Criteria for placement and algorithms for implementation are also discussed.

• Informal Description:

Sometimes commonly used public facilities (like a fire station or a shopping mall) are built where there is available land. Perhaps it would be better to build them in places which would maximize their use by the general public or facilitate their use by emergency personel. This module will address such questions and discuss possible methods of solution.

• Target Audience:

The intended target audience is high school seniors of average mathematical ability who are taking a course in discrete mathematics. However the module would certainly be appropriate for college students in a similar course.

• Prerequisites:

Students should be comfortable with the concept of graph and using a graph to model a physical situation. They should be familiar with basic matrix operations. Ability to use a TI-83 calculator would also be helpful.

• Mathematical Field:

Graph Theory (including weighted graphs) Elementary Matrix Operations

• Applications Areas:

The author’s intended application for this module is pedagogical. A student with a basic understanding of graphs and matrices can now see how these can be applied in a ”real world” setting.

• Mathematics Subject Classification:

Primary Classification: 90B80, 05C85

• Contact Information:

R. S. Tiberio Wellesley High School 50 Rice Street Wellesley, MA 02481 email: tiberio@tiac.net

• Other DIMACS modules related to this module: None at this time